(e) Express Laplace transforms of the following functions in terms of F(8)- (f(t)) 0 0
Answer #1


1,t2aBy second shifting theorem, By definition, ua(t) = u(-a) 0, t <a || L {t-a)u (t-a)) = e-u F(s), where LUC) = F (s) J